# Comprehensive Review of Deep Unfolding Techniques for Next-Generation Wireless Communication Systems

### v2

@@ Abstract
The application of machine learning in wireless communications has been extensively explored, with deep unfolding emerging as a powerful model-based technique. Deep unfolding enhances interpretability by transforming complex iterative algorithms into structured layers of deep neural networks (DNNs). This approach seamlessly integrates domain knowledge with deep learning (DL), leveraging the strengths of both methods to simplify complex signal processing tasks in communication systems. To provide a solid foundation, we first present a brief overview of DL and deep unfolding. We then explore the applications of deep unfolding in key areas, including signal detection, channel estimation, beamforming design, decoding for error-correcting codes, sensing and communication, power allocation, and security. Each section focuses on a specific task, highlighting its significance in emerging 6G technologies and reviewing recent advancements in deep unfolding-based solutions. Finally, we discuss the challenges associated with developing deep unfolding techniques and propose potential improvements to enhance their applicability across diverse wireless communication scenarios.
@@
Deep learning (DL), deep unfolding, model-driven machine learning, wireless communication signal processing, ISAC, IRS, THz, mmWave.
@@ I Introduction
The future mobile generations are expected to experience a surge in resource demand due to the exponential growth of connected devices. With the advent of 5G, billions of devices require seamless connectivity, intensifying competition for efficient spectrum allocation [1, 2, 3]. Each network has unique requirements and constraints, making resource allocation a complex challenge. Given the inherent limitations of available resources, problem formulation must prioritize flexibility, adaptability, and efficient utilization. Significant advancements in physical layer technologies have enhanced wireless communication systems, including massive multiple-input multiple-output (MIMO), intelligent reflecting surfaces (IRS), millimeter-wave (mmWave) communication, physical layer security, and non-orthogonal multiple access (NOMA). Additionally, orthogonal time-frequency space (OTFS) modulation is emerging as a promising alternative to orthogonal frequency-division multiplexing (OFDM), further improving spectral efficiency and resilience in next-generation networks.
In the communication system, traditional model-based analytical algorithms often employ iterative optimization techniques to achieve accurate solutions. Increasing the number of iterations improves performance as the algorithm converges closer to the optimal solution. Deriving an optimal solution in the form of a closed-form expression can be highly challenging in specific arbitrary communication scenarios. Hence, we resort to a sub-optimal solution in the traditional model-based method. Moreover, practical implementation on hardware imposes significant constraints on computational resources, latency, and power consumption, which limit the feasibility of deploying highly iterative methods. The classical iterative approach often relies on mathematically tractable models and algorithms for specific scenarios. Hence, adapting to dynamic and heterogeneous conditions, such as rapidly changing channels, interference, and diverse network configurations, can be challenging. Machine learning techniques offer a transformative approach by leveraging data-driven learning to extract patterns, adapt to new conditions, and approximate solutions efficiently. These methods can generalize across varying scenarios, learn optimal strategies from experience, and enable the design of intelligent communication systems that operate effectively in complex, real-world environments. These methods can potentially learn task-specific patterns and approximate solutions with significantly reduced computational complexity. Machine learning models can provide efficient and scalable alternatives to traditional iterative algorithms, enabling better performance under hardware limitations. This paradigm shift has opened new avenues for addressing complex problems in communication systems, bridging the gap between theoretical advancements and practical deployment.
@@
Deep Learning (DL), a machine learning subclass maps the relationship between input and output without explicitly modeling it in the architecture [4]. It involves the use of deep neural networks (DNN) with multiple hidden layers (hence “deep”) to analyze and interpret data. Numerous architectures have been proposed such as multilayer perceptron (MLP), convolutional NN (CNN), deep reinforcement learning (DRL), generative adversarial networks (GANs), etc. DNNs have been applied to various components of wireless communication, including detection, channel estimation, beamforming designs, and power allocation [5, 6, 7, 8, 9]. DL architectures exhibit several inherent limitations. These include complex training procedures and a black-box nature. The training process for these models requires extensive datasets, considerable time, and intricate fine-tuning. The complexity of implementing DNN architectures often constrains their applicability in real-world systems. Therefore, DL models are highly sensitive to the specific configurations of wireless systems, such as the number of antennas, users, and operating frequencies, requiring retraining or reimplementation whenever these parameters change. These challenges are not typically encountered in conventional optimization techniques, underscoring the potential advantages of using traditional algorithms [10, 11].
@@
Communication problems are inherently influenced by dynamic physical constraints that evolve in real-time. These constraints stem from factors such as the physical environment, hardware limitations, and the specific operational requirements of communication systems. Different communication scenarios impose varying demands on bandwidth, power consumption, and latency, which fluctuate with user density, data rates, and service types. Interpreting and addressing these constraints using conventional DNNs presents significant challenges. Traditional data-driven neural network architectures were developed primarily to process language and vision data, rendering them suboptimal for wireless communication tasks that involve channels, signals, and outputs governed by complex constraints. Although certain constraints can be handled through the application of specific activation functions, for example, employing the tanh function to enforce a constant modulus constraint in analog beamforming, standard activation functions generally lack the versatility to manage the intricate constraints characteristic of wireless communication systems. To mitigate these challenges, model-driven DL approaches facilitate the development of problem-specific network architectures and customized activation functions, optimized for the unique signal processing requirements of wireless communication applications [12].
Compared to the conventional fully-connected DL architectures, model-driven DL have fewer trainable parameters, implying less training time and computational complexity. Deep unfolding is an approach that falls into the category of model-driven DL [13]. The main motive behind developing a hybrid model (model-driven DL) for decision mapping, is to utilize the benefits of both optimization and data-driven DL. Model-driven DL is categorized into three frameworks as given in [13] : (1) learned optimizers [14, 15] (2) deep unfolding [16] (3) DNN-aided optimizers [17]. Here, we will focus on the second category i.e., deep unfolding. A deep unfolded network can be trained in an easier and faster fashion with smaller datasets than conventional DNNs. This is because deep unfolding incorporates domain knowledge from iterative algorithms into layers of DNNs, offering improved performance and interpretability [18]. The main idea in solving communication signal processing tasks lies in optimizing key hyperparameters of conventional algorithms based on data training. This includes elements such as step sizes in gradient-based methods, initial solutions, searching directions, criteria for early termination, and other relevant parameters. Deep unfolding uses suitable loss functions to train learnable parameters that are interpretable. Each iteration of the iterative algorithm is incorporated as a hidden layer of the DNN. These layers are then stacked to form a network, mimicking the optimization flow of algorithms. Furthermore, in many cases, the trainable parameters are independent of the system configurations, implying that the trained unfolding model can work for various scenarios.
@@ I-A Related Works
The growing application of deep unfolding in communication systems highlights the need for a comprehensive survey in this area. Wang et al. in [19] presented a review of DL and deep unfolding approaches for addressing physical layer signal processing challenges, including MIMO detection, channel estimation, and channel decoding. The authors of [20] provided an overview of deep unfolding applications in transceiver design, MIMO detection, and CSI estimation and feedback. Similarly, in [21], the authors highlighted the role of DL in communication, with a focus on DNNs, GANs, RNNs, and end-to-end learning. This work also explored deep unfolding for Bayesian optimal estimation, using MIMO detection and channel estimation as case studies. The authors in [22] focused primarily on MIMO detection while exploring its use in multiuser MIMO precoding and belief propagation-based channel decoding. A survey on model-based, DL, and hybrid approaches in the context of wireless communication networks is provided in [23]. In [16], Monga et al. conducted a survey of deep unfolding algorithms specifically implemented for applications in signal and image processing. The authors in [24] reviewed the implementation of deep unfolding in signal processing tasks. A brief review on deep unfolding in MIMO detection is presented in [17]. The authors in [13] have presented a general framework for optimization that leverages the benefit of deep unfolding with signal detection as one of the case studies. In [12], a chapter is dedicated to model-based machine learning for communication. The deep unfolding networks have been discussed for signal detection, channel estimation, and channel coding. Machine learning-based optimization framework is discussed in [25] with an emphasis on deep unfolding in specific design paradigms such as signal detection. A comparative study between purely data-driven and deep unfolding is conducted with complete emphasis on hybrid beamforming optimization in [26]. An overview of existing surveys on deep unfolding is shown in TABLE I.
@@ I-B Contributions
Existing surveys lack coverage of recent studies addressing critical 6G technologies, such as mmwave spectrum, new air interface that includes massive MIMO, IRS, next generation multiple access, and are often limited to specific application areas within the physical layer. Consequently, a comprehensive survey is needed to encompass the broad applications of deep unfolding techniques across various aspects of the physical layer in wireless communication. This paper provides an in-depth review of deep unfolding methods applied to the physical layer of wireless communication. The contributions are summarized below.
@@
    We provide a comprehensive overview of deep unfolding algorithms and their applications in physical layer wireless communication. Part A of the paper introduces the fundamentals of DL and deep unfolding to establish a foundation, outlining their general structures. We describe the DL architectures, including MLP, CNN, RNN, DRL, GNN, and GAN. Additionally, we detail a systematic, step-by-step approach for developing deep unfolding algorithms. Finally, we highlight the advantages and motivations for adopting hybrid approaches to tackle optimization tasks effectively.
@@
    We survey the application of deep unfolding in the areas of signal detection, channel estimation, precoder design, sensing and communication, decoding, power allocation, and physical layer security in Part B. Within each section, we present these areas with the key 6G technologies (like IRS, mmWave, etc.) used to aid readers to design frameworks that adapt to future technologies. Additionally, this survey explores the application of deep unfolding in integrated sensing and communication (ISAC).Also, we discuss the joint techniques in the physical layer in Part C. The different areas are joint detection and channel estimation, joint activity detection and channel estimation, joint channel estimation and precoder design, and joint detection and decoding.
@@
    We outline a set of challenges in developing deep unfolding models and propose future directions that could guide advancements in deep unfolding techniques to enhance their performance.
TABLE I: Existing Survey Papers on Deep Unfolding for Communication.
@@ I-C Paper Outline
  • Part A: We discuss the general structure of DNN and deep unfolding models. Section II covers the architecture of DNN and outlines the steps in designing a DNN. Section III presents the deep unfolding framework, which combines classical iterative algorithms with DNNs.
  • Part B: The applications of deep unfolding in different physical layer signal processing tasks are thoroughly reviewed. Specifically, Section IV explains signal detection with perfect knowledge of channel gains in the receiver, followed by channel estimation in Section V. Section VI covers the precoder design. Section VII reviews different deep unfolding approaches for ISAC. The decoding for error correction codes is discussed in Section VIII, followed by power allocation in Section IX and security aspects in Section X.
  • Part C : Section XI surveys the joint techniques in the physical layer. In this section, we discuss the application of deep unfolding in different joint techniques such as joint detection and channel estimation, joint activity detection and channel estimation, joint channel estimation and precoder design, and joint detection and decoding.
  • Part D : Section XII discusses the challenges in adopting the deep unfolding approach in wireless communication and the future scopes of research. Section XIII concludes the paper.
TABLE II: Acronyms and Corresponding Full Names.
Abbreviation Description Abbreviation Description
AMP Approximate message passing ADMM Alternating direction method of multipliers
BP Belief propagation BS Base station
BER Bit error rate CIR Channel impulse response
CNN Convolutional neural network CSI Channel state information
DL Deep learning DNN Deep neural network
DRL Deep reinforcement learning EM Expectation maximization
EP Expectation propagation FDD Frequency division duplexing
FFT Fast Fourier transform FDMA Frequency division multiple access
GAN Generative adversarial network GNN Graph neural network
HetNets Heterogeneous networks IRS Intelligent reflecting surface
IoT Internet of things ICI Inter-carrier interference
ISI Inter-symbol interference ISAC Integrated sensing and communication
ISTA Iterative shrinkage-thresholding algorithm JCEHB Joint channel estimation hybrid beamforming
JCDD Joint channel estimation detection and decoding LLR Log-likelihood ratio
LDPC Low-density parity check LPBN Long term passive beamforming
LMMSE Linear minimum mean square error LR Low resolution
LS Least squares LMS Least mean squares
LSTM Long short term memory MAP Maximum a posteriori
MDP Markov decision process MF Matched filtering
MO Manifold optimization MM Majorization minimization
MS Mobile station MMSE Minimum mean square error
MIMO Multiple input multiple output ML Maximum likelihood
MPD Message passing detector MSE Mean square error
MSER Minimum symbol error rate mMTC Massive machine type communication
NN Neural network NOMA Non-orthogonal multiple access
OAMP Orthogonal approximate message passing OFDM Orthogonal frequency division multiplexing
OMP Orthogonal matching pursuit OTFS Orthogonal time frequency space
PGA Projected gradient ascent PGD Projected gradient descent
RF Radio frequency RNN Recurrent neural network
RL Reinforcement learning RSMA Rate splitting multiple access
SSCA Stochastic successive convex approximation SABN Short term active beamforming network
SBRNN Sliding bidirectional recurrent neural network SCDMA Sparse code division multiple access
SE Spectral efficiency SGD Stochastic gradient descent
SIC Successive interfernce cancellation SISO Single-input single-output
SIMO Single-input multiple-output SINR Signal-to-interference and noise ratio
SNR Signal-to-noise ratio TAS Transmit antenna selection
TDD Time division duplexing THz Tera Hertz
UAV Unmanned aerial vehicle URLLC Ultra reliable low latency communication
WMMSE Weighted minimum mean square error ZF Zero forcing
@@ I-D Notation and Acronyms
II Deep Neural Network: General Structure
Artificial intelligence originated with the famous Turing test in the 1950s, marking a breakthrough in machine thinking capability [28]. Over the decades, around the 1990s, machine learning formed a subclass under artificial intelligence. Machine learning requires manual feature extraction of structured data. Most machine learning algorithms can be categorized into supervised, unsupervised, and reinforcement learning [4].
Figure 2: Feedforward NN (MLP) architecture.
Supervised learning aims to learn how to map inputs to outputs to predict labels for new data. In unsupervised learning, the dataset includes many features that are not labeled. The algorithm tries to identify hidden patterns within the data without the guidance of labeled examples. The third category, reinforcement learning, consists of an agent that learns to make decisions based on the environment. Since there is no predefined dataset, the agent interacts with the environment and learns from the feedback it receives for every action [4],[29]. The output of a machine learning algorithm can be of two types: regression and classification.
In the 2010s, DL emerged as a crucial technology in designing intelligent networks with unstructured data [30]. Researchers have adopted DL techniques to solve complex problems. It has been widely used in computer vision, natural language processing, and image processing. This success has inspired researchers to use DL in the physical layer of wireless communication, especially in various signal-processing tasks.
An NN consists of interconnected layers of neurons that work together to learn patterns from the given data. Different types of artificial NN architectures have been developed for various learning tasks, such as feed-forward NNs, CNN, recurrent NN (RNN), graph NN (GNN), etc. Fig. 2 shows the feedforward NN architecture with $K$ ($K=2$ here) hidden layers and $N_{k}$ neurons in layer $k$, for all $k=0,\dots,K$. The architecture has an input layer, hidden layers, and an output layer.
We now discuss the steps in a feedforward NN:
  1. 1.
    The network begins with an input vector
    ${\bf{s}}=[s_{1},s_{2},\dots,s_{n}]$, at the input layer where $n$ is the number of input features.
  2. 2.
    For each neuron in a layer, the weighted sum of inputs is computed:
    $z^{(k)}_{j}=\sum_{i=1}^{n}w_{ji}^{(k)}s_{j}+b_{j}^{(k)},$
    where $w_{ji}^{(k)}$ is the weight connecting neuron $i$ to neuron $j$ in the $k^{\text{th}}$ layer and $b_{j}^{(k)}$ is the bias for neuron $j$ at the $k^{\text{th}}$ layer.
  3. 3.
    A non-linear activation function $\sigma(\cdot)$ is applied:
    $a_{j}^{(k)}=\sigma(z^{(k)}_{j}),$
    where $a_{j}^{(k)}$ is the activation output of the neuron $j$ in the layer $k$. Common activation functions include ReLU, sigmoid, tanh, softmax etc. [4].
  4. 4.
    At the output layer, after applying the weighted sum and activation process, the output is given by:
    $\hat{\bf{y}}=\sigma(\mathbf{W}^{K}\mathbf{s}^{K}+\mathbf{b}^{K})$
    where ${\bf{W}}$ is the weight matrix at the output layer and $\bf{b}$ is the bias vector at the output layer. The activation function, ($\sigma(\cdot)$) for the output often depends on the task, whether it is a classification or regression task.
  5. 5.
    The loss is computed to measure the network’s performance:
    $\mathcal{L}(\hat{\bf{y}},\bf{y})$
    where $\hat{\bf{y}}$ is the predicted output, and $\bf{y}$ is the true output. The common loss functions are mean squared error (MSE), binary cross entropy, categorical cross entropy, etc.
To train the NN, backpropagation is used by adjusting the parameters of the network, i.e., the weight and biases to minimize the loss. The parameters are updated using an optimization algorithm like stochastic gradient descent (SGD).
On the other hand, CNN is designed to process grid-like data by leveraging spatial structure. It uses layers of convolutional filters to extract features like edges followed by pooling layers to reduce spatial dimensions and retain key information [31]. The extracted features are passed through fully connected layers. One of the early applications of CNN was modulation classification based on signal constellation diagrams; the authors have used the CNN-based DL method [32]. They have developed and analyzed methods to represent modulated signals in data formats with grid-like topologies to conform with the CNN architecture.
RNN is another type of NN designed explicitly for processing sequential data. In RNN, the output from the previous step is fed as an input to the current step, incorporating sequential dependencies between the nodes. A recurrent neuron performs this function by maintaining a hidden state. Long short-term memory (LSTM) is one version of RNN [4]. The LSTM network was adopted into the NOMA framework in [33] to learn the channel characteristics of the NOMA system.
DRL integrates RL with DL to develop intelligent agents capable of making sequential decisions in complex environments. In DRL, an agent interacts with its environment to maximize cumulative rewards, learning optimal policies through exploration and exploitation. DNNs are employed to approximate value functions or policies, allowing DRL to efficiently handle high-dimensional state and action spaces. Key techniques such as Deep Q-Networks (DQN), policy gradient methods, and actor-critic frameworks enable agents to improve their performance iteratively through trial and error, guided by reward signals [4, 34]. In the domain of communications and networking, DRL has emerged as a powerful tool for addressing challenges such as resource allocation, spectrum management, and network optimization. Modern network architectures, including the Internet of Things (IoT), heterogeneous networks (HetNets), and unmanned aerial vehicle (UAV) networks, exhibit decentralized, dynamic, and autonomous characteristics, making them well-suited for reinforcement learning-based solutions [35].
The concept of using autoencoders to design end-to-end communication systems has been widely studied [36]. End-to-end learning trains a single model to map raw input directly to the desired output. It avoids intermediate steps or manual feature engineering. Autoencoders are neural networks that optimize both the transmitter and receiver together in this framework. In communication systems, the goal is to design the transmitter and receiver jointly for a given channel model. This helps achieve better performance, such as faster convergence and lower bit error rates [37].
GNNs are specialized networks designed for graph-structured data, where nodes are connected by edges [38]. Unlike traditional neural networks, GNNs work on non-Euclidean data and capture relationships between graph nodes. They learn rich representations of nodes, edges, or the whole graph by aggregating and transforming features from neighboring nodes. This ability to use graph structure helps GNNs perform better in tasks that rely on relational data.
In 2020, Goodfellow et al. proposed a new AI algorithm for generative modeling [39]. GANs use two neural networks: a generator and a discriminator. The generator creates realistic outputs, while the discriminator identifies if they are real or fake. By competing and learning together, GANs can produce high-quality outputs that are often indistinguishable from real data.
For further reading on application of deep learning in physical layer of communication, we invite readers to refer [40, 41, 42, 27].
III Deep Unfolding: General Structure
Deep unfolding, also known as algorithm unrolling [16], is a typical model-driven DL method [10],[20],[43, 44, 45, 46, 47, 48]. It has been frequently considered for various tasks such as signal detection, design of beamforming matrices, and many more optimization tasks in wireless communication. Owing to reduced training time and dataset, deep unfolding is one promising alternate approach to purely model-driven or data-driven.
Figure 3: Schematics of data-driven, classical and deep-unfolding approaches.
Deep unfolding integrates model-based approaches and domain knowledge into the design of DNN architectures. The training is carried out in an end-to-end manner. In deep unfolding, an iterative algorithm is mapped to DNN layers. Then, DL techniques such as SGD are used to learn the parameters. Fig. 3 shows a schematic diagram comparing the classical approach, data-driven approach, and the deep unfolding approach. Furthermore, Fig. 4 depicts the deep unfolding approach in greater details.

### v3

@@ Abstract
The massive surge in device connectivity demands higher data rates, increased capacity with low latency and high throughput. Hence, to provide ultra-reliable, low-latency communication with ubiquitous connectivity for Internet-of-Things (IoT) devices, next-generation wireless communication leverages the incorporation of machine learning tools. However, standard data-driven models often need large datasets and lack interpretability. To overcome this, model-driven deep learning (DL) approaches combine domain knowledge with learning to improve accuracy and efficiency. Deep unfolding is a model-driven method that turns iterative algorithms into deep neural network (DNN) layers. It keeps the structure of traditional algorithms while allowing end-to-end learning. This makes deep unfolding both interpretable and effective for solving complex signal processing problems in wireless systems. We first present a brief overview of the general architecture of deep unfolding to provide a solid foundation. We also provide an example to outline the steps involved in unfolding a conventional iterative algorithm. We then explore the application of deep unfolding in key areas, including signal detection, channel estimation, beamforming design, decoding for error-correcting codes, integrated sensing and communication, power allocation, and physical layer security. Each section focuses on a specific task, highlighting its significance in emerging 6G technologies and reviewing recent advancements in deep unfolding-based solutions. Finally, we discuss the challenges associated with developing deep unfolding techniques and propose potential improvements to enhance their applicability across diverse wireless communication scenarios.
@@
Deep learning, Deep unfolding, Model-driven machine learning, Wireless communication signal processing, Integrated sensing and communication (ISAC), Intelligent reflecting surfaces (IRS), Terahertz, IoT.
@@ I Introduction
The future mobile generations are expected to experience a surge in resource demand due to the exponential growth of connected devices. With the advent of 5G, billions of devices require seamless connectivity, intensifying competition for efficient spectrum allocation [1, 2, 3]. Each network has unique requirements and constraints, making resource allocation a complex challenge. Given the inherent limitations of available resources, problem formulation must prioritize flexibility, adaptability, and efficient utilization. Significant advancements have taken place in wireless communication due to various exciting physical layer technologies, including massive multiple-input multiple-output (MIMO), Internet-of-Things (IoT), integrated sensing and communication (ISAC), intelligent reflecting surfaces (IRS), millimeter-wave (mmWave) communication, physical layer security, and next-generation multiple access such as nonorthogonal multiple access (NOMA). Additionally, orthogonal time-frequency space (OTFS) modulation is emerging as a promising alternative to orthogonal frequency-division multiplexing (OFDM), further improving spectral efficiency (SE) and resilience in next-generation networks.
@@
Deep learning (DL), a machine learning subclass, maps the relationship between input and output without explicitly modeling it in the architecture [4]. It involves the use of deep neural network (DNN) with multiple hidden layers (hence “deep”) to analyze and interpret data. Numerous architectures have been proposed such as multilayer perceptron (MLP), convolutional neural network (CNN), deep reinforcement learning (DRL), generative adversarial networks (GANs), etc. DNNs have been applied to various components of wireless communication, including detection, channel estimation, beamforming designs, and power allocation [5, 6, 7, 8, 9]. However, DL architectures exhibit several inherent limitations. These include complex training procedures and a black-box nature of algorithms. The training process for these models requires extensive datasets, considerable time, and intricate fine-tuning. The complexity of implementing DNN architectures often constrains their applicability in real-world systems. Therefore, DL models are highly sensitive to the specific configurations of wireless systems, such as the number of antennas, users, and operating frequencies, which require retraining or reimplementation whenever these parameters change. These challenges are not typically encountered in conventional optimization techniques, underscoring the potential advantages of using classical algorithms [10, 11].
@@
Communication problems are inherently influenced by dynamic physical constraints that evolve in real-time. These constraints stem from factors such as the physical environment, hardware limitations, and the specific operational requirements of communication systems. Different communication scenarios impose varying demands on bandwidth, power consumption, and latency, which fluctuate with user density, data rates, and service types. Interpreting and addressing these constraints using DNNs presents significant challenges. Data-driven neural network (NN) architectures were developed primarily to process language and vision data, rendering them suboptimal for wireless communication tasks that involve channels, signals, and outputs governed by complex constraints. Although certain constraints can be handled through the application of specific activation functions, for example, employing the $tanh$ function to enforce a constant modulus constraint in analog beamforming, standard activation functions generally lack the versatility to manage the intricate constraints characteristic of wireless communication systems. To mitigate these challenges, model-driven DL approaches facilitate the development of problem-specific network architectures and customized activation functions, optimized for the unique signal processing requirements of wireless communication [12].
Compared to the conventional fully-connected DL architectures, model-driven DL have fewer trainable parameters, implying less training time and computational complexity. Deep unfolding is an approach that falls into the category of model-driven DL [13]. There are several types of model-driven DL approaches, including deep unfolding [14],[15], physics informed NN [16, 17], learned optimizers [18, 19], etc. A deep unfolded network can be trained in an easier and faster fashion with smaller datasets than DNNs. This is because deep unfolding incorporates domain knowledge from iterative algorithms into layers of DNNs, offering improved performance and interpretability [20].
TABLE I: Acronyms and corresponding full names.
Abbreviation Description
AO Alternate optimization
AP Access point
AMP Approximate message passing
ADC Analog-to-digital converters
ADMM Alternating direction method of multipliers
BP Belief propagation
BS Base station
BER Bit error rate
CD Coordinate descent
CNN Convolutional neural network
CSI Channel state information
CGD Conjugate gradient descent
DL Deep learning
DNN Deep neural network
DRL Deep reinforcement learning
EM Expectation maximization
EP Expectation propagation
FDD Frequency division duplexing
GAN Generative adversarial network
GNN Graph neural network
HBF Hybrid beamforming
IRS Intelligent reflecting surface
IoT Internet-of-things
ISI Inter-symbol interference
ISAC Integrated sensing and communication
ISTA Iterative shrinkage-thresholding algorithm
LR Low-resolution
LS Least squares
LLR Log-likelihood ratio
LDPC Low-density parity check
LMMSE Linear minimum mean square error
MF Matched filter
ML Maximum likelihood
MO Manifold optimization
MU Multi-user
MAP Maximum a posteriori
MPD Message passing detector
MSE Mean square error
MMSE Minimum mean square error
MIMO Multiple input multiple output
MISO Multiple input single output
MSER Minimum symbol error rate
mMTC Massive machine type communication
mmWave millimeter-wave
NN Neural network
NOMA Non-orthogonal multiple access
NMSE Normalized mean square error
OAMP Orthogonal approximate message passing
OFDM Orthogonal frequency division multiplexing
OTFS Orthogonal time frequency space
PGA Projected gradient ascent
PGD Projected gradient descent
QAM Quadrature amplitude modulation
QoS Quality of service
QPSK Quadrature phase shift keying
RF Radio frequency
RNN Recurrent neural network
RSMA Rate splitting multiple access
SD Sphere decoder
SE Spectral efficiency
SER Symbol error rate
SCA Successive convex approximation
SGD Stochastic gradient descent
SIC Successive interference cancellation
SNR Signal-to-noise ratio
SSCA Stochastic successive convex approximation
SINR Signal-to-interference and noise ratio
THz Terahertz
TPGD Trainable projected gradient descent
WMMSE Weighted minimum mean square error
ZF Zero forcing
@@ I-A Related Works
The growing adoption of deep unfolding in communication systems has led to several surveys covering its diverse applications. Wang et al. [21] reviewed DL and deep unfolding methods for physical layer tasks such as MIMO detection, channel estimation, and decoding. The authors of [22] discussed its use in transceiver design and channel state information (CSI) estimation. In [23], deep learning methods including DNNs, GANs, and recurrent NN (RNN) were reviewed, with deep unfolding applied to Bayesian estimation. The authors of [15] focused on MIMO detection, multiuser precoding, and belief propagation. A broader comparison of model-based, DL, and hybrid methods is found in [24]. Monga et al. [14] focused on algorithm unrolling for signal/image processing. Additional works [25, 13] explored signal processing and optimization. Eldar et al. [12] and Shi et al. [26] addressed DL-based communication frameworks. In [27], the authors highlight the advantages of deep unfolding for physical receiver design, with particular emphasis on its applications in signal detection and channel estimation. A comparison between data-driven and unfolding-based beamforming appears in [28] and Table II.
@@ I-B Contributions
Existing surveys lack coverage of recent studies addressing critical 6G technologies, such as mmWave spectrum, new air interface that includes massive MIMO, IRS, next-generation multiple access, and are often limited to specific application areas. Consequently, a comprehensive survey is needed to encompass the broad applications of deep unfolding techniques across various aspects of the physical layer in wireless communication. This paper provides an in-depth review of deep unfolding methods applied to the physical layer of wireless communication. The contributions are summarized below:
@@
    This paper presents a systematic, step-by-step approach for developing deep unfolding algorithms, with a detailed tutorial on the alternating direction method of multipliers (ADMM)-based MIMO detection. The tutorial offers insights into the ADMM unfolding process, identifies trainable parameters, and demonstrates the design of the resulting unfolded network.
@@
    We survey the application of deep unfolding across various domains, including signal detection, channel estimation, precoder design, ISAC, decoding, power allocation, and physical layer security. Each section highlights the integration of deep unfolding with key 6G technologies such as IRS, mmWave, and IoT, providing readers with guidance on designing frameworks adaptable to future wireless systems.
@@
    In addition, each section concludes with a comparative analysis table that evaluates prominent unfolded networks across key metrics such as computational complexity, generalization capability, and convergence behavior. These tables highlight the practical trade-offs and deployment boundaries associated with each architecture. Furthermore, we discuss the key takeaways to provide readers with deeper insights into the practical applicability and design considerations of deep unfolding in physical layer communication.
  4. 4.
    This paper concludes the survey by outlining a set of challenges in developing deep unfolding models and propose future directions that could guide advancements in deep unfolding techniques to enhance their performance.
TABLE II: Existing survey papers on deep unfolding for communication.
@@ 11 Machine Learning and Wireless Communications [12] 2022 $\checkmark$ $\checkmark$ $\checkmark$
14 Model and Data Dual-Driven: A Promising Receiver Design for 6G Wireless Communications [27] 2025 $\checkmark$ $\checkmark$
@@ I-C Paper Outline
  • Part A: Section II presents the deep unfolding framework, which combines classical iterative algorithms with DNNs. The understanding of deep unfolding is exemplified via ADMM based MIMO detection. A step-by-step approach is presented for developing the unfolding of ADMM specifically for MIMO detection.
  • Part B: This part explores deep unfolding applications in physical layer tasks: signal detection (Section III), channel estimation (Section IV), precoder design (Section V), ISAC (Section VI), decoding for error correcting code (Section VII), power allocation (Section VIII), and security (Section IX).
  • Part C : Section X discusses the challenges in adopting the deep unfolding approach in wireless communication and the future scope of research. Section XI concludes the paper.
@@ I-D Notation and Acronyms
II Deep Unfolding: General Architecture and Fundamentals
Deep unfolding, or algorithm unrolling [14], bridges model-based iterative algorithms and data-driven DNNs [10], [22], [31], [32], [33], [34], [35], [36]. It maps each iteration of an iterative algorithm to a DNN layer, enabling end-to-end learning while retaining interpretability. In deep unfolding, the learnable parameters include model-specific variables and regularization coefficients, optimized layer by layer. Fig. 2 illustrates the process of unfolding an iterative algorithm. In general, for any iterative algorithm, the update equation for each iteration is given as:
${\bf{s}}^{l+1}=h({\bf{s}}^{l};{\boldsymbol{\theta}}^{l})$ (1)
where $\left\{{\bf{s}}^{l}\right\}_{l=1}^{L}\in{\mathbb{R}}^{n}$ denotes the sequence of variable vector updated across $L$ layers of the algorithm. At each layer, the update is performed by executing $h(\cdot;\cdot)$. Repeated application of this update rule over $L$ layers, yields the unfolded network architecture, $h(\cdot;{\boldsymbol{\theta}}^{l})_{l=1}^{L}$. The trainable parameters, $\left\{\boldsymbol{\theta}^{l}\right\}_{l=1}^{L}$ (shown in red fonts in Fig. 2) includes the model parameters and the step-size for the regularization coefficients of the algorithm [26].